# Temporal Inductive Path Neural Network for Temporal Knowledge Graph Reasoning

### 08 Dec 2023 v2

@@
Abstract.
Temporal Knowledge Graph (TKG) is an extension of traditional Knowledge Graph (KG) that incorporates the dimension of time. Reasoning on TKGs is a crucial task that aims to predict future facts based on historical occurrences. The key challenge lies in uncovering structural dependencies within historical subgraphs and temporal patterns. Most existing approaches model TKGs relying on entity modeling, as nodes in the graph play a crucial role in knowledge representation. However, the real-world scenario often involves an extensive number of entities, with new entities emerging over time. This makes it challenging for entity-dependent methods to cope with extensive volumes of entities, and effectively handling newly emerging entities also becomes a significant challenge. Therefore, we propose Temporal Inductive Path Neural Network (TiPNN), which models historical information in an entity-independent perspective. Specifically, TiPNN adopts a unified graph, namely history temporal graph, to comprehensively capture and encapsulate information from history. Subsequently, we utilize the defined query-aware temporal paths to model historical path information related to queries on history temporal graph for the reasoning. Extensive experiments illustrate that the proposed model not only attains significant performance enhancements but also handles inductive settings, while additionally facilitating the provision of reasoning evidence through history temporal graphs.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Temporal reasoning
1. Introduction
@@
Knowledge Graphs (KGs) are powerful representations of structured information that capture a vast array of real-world facts. They consist of nodes, which represent entities such as people, places, and events, and edges that define the relationships between these entities. These relationships are typically represented as triples in the form of $(s,r,o)$, comprising a subject, a relation, and an object, such as (Tesla, Founded by, Elon Musk). Temporal Knowledge Graphs (TKGs) is an extension of a traditional KGs that incorporates the dimension of time. In a TKG, these facts are annotated with timestamps or time intervals to indicate when they are or were valid. Different from KGs, each fact in a TKG is represented as a quadruple, adding a timestamp feature, which can be denoted as $(s,r,o,t)$, including a subject, a relation, an object, and a timestamp. For example, (Albert Einstein, Born in, Germany, 03/14/1879). The timestamp indicates the specific date or time when this event occurred, providing temporal information to the graph. Such unique representation empowers TKGs to capture the dynamics of multi-relational graphs over time, exhibiting temporal variations. As a result, TKGs have found widespread applications in diverse domains, such as social network analysis (Wang et al., 2022)(Gervais, 2017), and event prediction (Jin et al., 2019)(Bai et al., 2021)(Wang et al., 2020) (e.g. disease outbreaks, natural disasters, and political shifts, etc.) These applications leverage the temporal information embedded in TKGs to make informed decisions and predictions based on historical knowledge.
@@
Among various tasks, TKG reasoning is one emerging research topic that focuses on inferring missing facts based on known ones. Basically, reasoning over TKG consists two primary settings: interpolation and extrapolation (Jin et al., 2020). Given the timestamp scope $[0,T]$ in a TKG, where $T$ represents the maximum timestamp, in the interpolation setting, the reasoning process is focused on the time range from 00 to $T$, i.e., completing missing facts within the range of past timestamps based on the given known data (Goel et al., 2020). On the other hand, in the extrapolation setting, the reasoning timestamps extend beyond $T$, focusing on predicting the facts in the future based on the observed historical knowledge encoded in the TKG (Li et al., 2021)(Dong et al., 2023)(Zhu et al., 2020). This study particularly emphasizes the extrapolation setting that infers future missing facts.
@@
Existing literature on TKG reasoning mainly focuses on exploring temporal features and structural information to capture complicated variations and logical connections between entities and relations. For example, RE-NET (Jin et al., 2020) and RE-GCN (Li et al., 2021) are representative works that leverage RNN-based and GNN-based approaches to learn both temporal dependencies and structural features from historical sequences in TKGs, serving as the foundation for predicting future facts. xERTE (Han et al., 2020a) utilizes the entity connection patterns from different temporal subgraphs in TKGs to construct an inference graph with time-aware features, enabling graph reasoning. TITer (Sun et al., 2021a) adopts a different approach by fusing historical subgraph sequences into a unified graph through the addition of extra edges. It employs a reinforcement learning framework and designs a reward function to guide the model learning. These methods rely on entity modeling as nodes in the graph play a crucial role in knowledge representation, and entity modeling is indispensable, as entities serve as important hubs and key components in the graph.
@@
However, the real-world scenario often involves an extensive number of entities, with new entities emerging over time. This makes it challenging for entity-dependent methods to cope with extensive volumes of entities, and effectively handling newly emerging entities also becomes a significant challenge. To address this, we previously introduced an innovative entity-independent modeling approach for temporal knowledge graphs with an evolutionary perspective in our previous work (Dong et al., 2023), referred to as DaeMon. The entity-independent modeling approach brings several benefits: (i) modeling on graph representation is independent of entities, maintaining stable efficiency even for large-scale datasets with numerous entities. In other words, the memory occupation is independent of the number of entity categories. (ii) the entity-independent nature makes it insensitive to entity-specific features. Hence, when new entities emerge in future timestamps, the model can still handle these unseen nodes, showcasing robust generalization and reasoning capabilities. In essence, DaeMon focuses on learning query-aware latent path representations, incorporating a passing strategy between adjacent timestamps to control the path information flow toward later ones without explicitly modeling individual entities. Specifically, DaeMon introduces the concept of temporal paths to represent the logical paths between a query’s head entity and candidate tail entities within the sequence of historical subgraphs. It utilizes path memory units to store aggregated representations of temporal paths along the timeline. The model further employs a path aggregation unit to capture local path information within each subgraph. Finally, DaeMon adopts a memory passing strategy to facilitate cross-temporal information propagation.
@@
Here, we provide a generalized overview of inference based on temporal paths modeling, as illustrated in Figure 1. For a given inference query, it starts from this query and collects the path between the head node and tail node in the historical context. Then, it utilizes a query-aware temporal method to represent the aggregation of temporal paths with temporal information for reasoning. For example, in Figure 1, the red triangular nodes and red circular nodes represent the head entity and candidate tail entity of the query at timestamp $T+1$, respectively. This method, based on temporal paths, iteratively learns the path information between the red triangular and red circular nodes within the historical context. The aggregation of path information corresponding to the query is then utilized to infer the possibility of potential future interactions. Besides, during the learning of path information, the method does not focus on which intermediary nodes (i.e., the blue nodes in Figure 1) are present in the path, but rather employs the relation representations present in the path to express it, showcasing an entity-independent characteristic.
Figure 1. A generalized overview of inference based on temporal paths modeling.
@@
While DaeMon benefits from this entity-independent modeling approach, it still exhibits certain limitations. First, during the learning of historical information, DaeMon independently performs graph learning on local subgraphs and then connects them through memory passing strategy to obtain the final query-aware path features for future fact reasoning. This process requires graph learning on each subgraph separately, leading to difficulties in modeling complex temporal characteristics and a significant increase in time complexity as the historical length grows. Second, the memory passing strategy is adopted to model temporal features in historical sequences, controlling the information flow from one timestamp to the next to achieve temporal evolution modeling. However, this temporal modeling approach is unidirectional, which not only gives rise to the long-term dependency problem but also neglects relative temporal features between timestamps. Third, DaeMon defines temporal paths as virtually connected paths across the sequence of subgraphs. However, in practice, it does not directly learn the representation of these virtual paths. Instead, it independently learns several edges on each timestamp’s subgraph and then obtains the path feature by fusing the edge information through memory passing strategy. The independent learning approach without real physical paths creates challenges in providing relevant reasoning evidence during the inference stage, which makes it difficult to present interpretable explanations to users, hindering their understanding of the reasoning process.
@@
To overcome the above challenges, we propose a model, Temporal Inductive Path Neural Network (TiPNN). Specifically, we introduce the concept of history temporal graphs as a replacement for the original historical subgraph sequences to mitigate the complexity of learning historical information independently on multiple subgraphs. It allows us to model historical information more efficiently by combining it into a single unified graph, thereby enabling the discovery of complex temporal characteristics and reducing the computational burden during model inference. Furthermore, we redefine the notion of temporal paths as a logical semantics path between the query entity and candidate entities in history temporal graph and design query-aware temporal path processing to model the path feature on the single unified graph in the entity-independent manner for the final reasoning. This enhancement enables the model to comprehensively learn both the temporal and structural features present in the history temporal graph. The adoption of the entity-independent modeling approach allows the model to effectively handle the inductive setting on the history temporal graph. By jointly modeling the information within and between different timestamps in a single unified graph, it can capture a more intuitive and holistic view of the temporal reasoning logic, thus facilitating user understanding of the model’s reasoning process and can even provide interpretable explanations for the reasoning results, meanwhile providing new insights for this task.
In essence, this work revolves around capturing query-aware temporal path features on the history temporal graphs to enable the inference of missing future facts. Our main contributions are as follows:
@@
  • We define the concept of temporal paths. Subsequently, we propose the Temporal Inductive Path Neural Network (TiPNN) for the task of temporal extrapolated reasoning over TKGs. Our approach models the temporal path features corresponding to reasoning queries by considering temporal information within and between different timestamps, along with the compositional logical rules of historical subgraph sequences.
@@
  • Extensive experiments demonstrate that our model outperforms the existing state-of-the-art results. We analyze the impact and effectiveness of individual components in TiPNN. Additionally, we construct datasets and conduct validation experiments for inductive reasoning tasks on temporal knowledge graphs. Furthermore, we present reasoning evidence demonstrations and analytical experiments based on history temporal graph inference.
2. Related Work
2.1. Knowledge Graph Reasoning
In recent years, knowledge graph representation learning has witnessed significant developments, aiming to embed entities and relations into continuous vector spaces while capturing their semantics (Bordes et al., 2013) (Sun et al., 2019) (Li et al., 2022a). These methods can be broadly categorized into three categories: translation-based methods, semantic matching methods, and neural network-based methods. Translation-based methods, such as TransE (Bordes et al., 2013). It considers relations as translations from subject entities to object entities in the vector space. Building upon TransE, several improved methods have been proposed, including TransH (Wang et al., 2014), TransR (Lin et al., 2015), TransD (Ji et al., 2015) and TransG (Xiao et al., 2015), which introduce various strategies to enhance the modeling of entity-relation interactions. For semantic matching methods, RESCAL (Nickel et al., 2011) proposes a tensor-based relational learning approach capable of collective learning. DistMult (Yang et al., 2014) simplifies RESCAL using diagonal matrices for efficiency, while HoIE (Nickel et al., 2016) and ComplEx (Trouillon et al., 2016a) extend the representation capacity by incorporating higher-order interactions and complex-valued embeddings (Liu et al., 2020), respectively. Neural network-based methods have also gained attention, with approaches like GCN (Welling and Kipf, 2016), R-GCN (Schlichtkrull et al., 2017), WGCN (Shang et al., 2019), VR-GCN (Ye et al., 2019), and CompGCN (Vashishth et al., 2019), which integrate content and structural features within the graph, allowing for joint embedding of nodes and relations in a relational graph (Ning et al., 2021) (Qiao et al., 2020). These methods effectively capture complex patterns and structural dependencies in KGs, pushing the boundary of KG representation learning. Despite the successes of existing methods in reasoning with static KGs, they fall short when it comes to predicting temporal facts due to the lack of temporal modeling.
2.2. Temporal Knowledge Graph Reasoning
For interpolation reasoning, the objective is to infer missing facts from the past within observed data (Dasgupta et al., 2018) (Esteban et al., 2016) (Garcia-Duran et al., 2018). TTransE (Leblay and Chekol, 2018) extends the TransE (Bordes et al., 2013) by incorporating relations and timestamps as translation parameters between entities. TA-TransE (García-Durán et al., 2018) integrates the timestamps corresponding to fact occurrences into the relation representations, while HyTE (Dasgupta et al., 2018) associates timestamps with corresponding hyperplanes. Additionally, based on ComplEx (Trouillon et al., 2016a), TNTComplEx (Lacroix et al., 2020) adopts a unique approach where the TKGs are treated as a 4th-order tensor, and the representations are learned through canonical decomposition. However, these methods are not effectively applicable for predicting future facts (Goel et al., 2020)(Han et al., 2020b) (Leblay and Chekol, 2018) (Sadeghian et al., 2016).
In contrast, the extrapolation setting, which this work focuses on, aims to predict facts in future timestamps based on historical TKG sequences. Know-Evolve (Trivedi et al., 2017) uses the temporal point process to represent facts in the continuous time domain. However, it falls short of capturing long-term dependencies. Similarly, DyREP (Trivedi et al., 2019) posits representation learning as a latent mediation process and captures the changes of the observed processes. CyGNet (Zhu et al., 2021a) utilizes a copy-generation mechanism that collects frequently repeated events with the same subject entities and relations to the query for inferring. RE-NET (Jin et al., 2020) conducts GRU and GCN to capture the temporal and structural dependencies in a sequential manner. xERTE (Han et al., 2020a) offers an explainable model for their predictions by searching in the subgraph with attentive propagation. RE-GCN (Li et al., 2021) considers both structural dependencies and static properties of entities at the same time, modeling in evolving manner. TANGO (Han et al., 2021) designs neural ordinary differential equations to represent and model TKGs for continuous-time reasoning. In addition, reinforcement learning methods are used to uncover the most related facts to answer the query, such as TITer (Sun et al., 2021a). Most existing methods are primarily transductive in nature, only focusing on modeling the graph features based on entities while overlooking the intrinsic temporal logical rules within TKG reasoning. As a result, there is still a need for more efficient methods that can effectively manage the robustness and scalability requirements in real-world situations.
3. Preliminaries
Table 1. Notations and Descriptions.
@@ 3. Preliminaries
3.1. Background of Temporal Knowledge Graph
@@
3.2. Formulation of Reasoning Task
@@ 4.1. Model Overview
4.2. History Temporal Graph Construction
@@
Figure 2. Example of constructing the history temporal graph $\hat{G}_{t-2:t}$ with $m=3$. For illustrative purposes, we use an undirected graph to demonstrate the construction method and omit the relationship types of edges in the historical subgraph sequence and the constructed history temporal graph. The time labels attached to the edges in the history temporal graph represent the timestamps of the corresponding edges in the historical subgraph sequence.
@@
4.3. Temporal Path Formulation
@@
Definition 4.1 (Temporal Path).
is a logical path that aggregates the semantics of all connected paths in history temporal graph between a head entity and a tail entity through arithmetic logical operations.
The aggregated temporal path can be used to simultaneously represent the temporal information and comprehensive path information from the head entity to the tail entity. This enables us to obtain a comprehensive query-aware paired representation for the future facts reasoning task by aggregating the information from various paths, which is from query subject node to object candidate nodes in history temporal graph. Note that it is essential to consider the relation type information from the query while learning paired representations. In other words, for a query $(s,r,?,t+1)$, we should also embed the relation type $r$ into the paired representation (i.e., path representation) as well. This ensures that the relation type is incorporated into the learning process, enhancing the overall modeling of the query and enabling more accurate and context-aware inference of missing facts, and the detail will be discussed in Section 4.4.
We denote the representation of temporal path in history temporal graph $\hat{G}_{\prec t+1}$ as $\textbf{H}^{\prec t+1}_{(s,r)\rightarrow\mathcal{V}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ corresponding to the query $(s,r,?)$ at future timestamp $t+1$, where $|\mathcal{V}|$ is the cardinality of object candidates set and $d$ is the dimension of temporal path embedding. That is, $\textbf{H}^{\prec t+1}_{(s,r)\rightarrow\mathcal{V}}$ describes the representations of temporal paths from query subject entity $s$ to all object candidate entities in $\mathcal{V}$, considering the specific query $(s,r,?)$. Specifically, each item in $\textbf{H}^{\prec t+1}_{(s,r)\rightarrow\mathcal{V}}$, which is denoted as $\textbf{h}^{\prec t+1}_{(s,r)\rightarrow o}\in\mathbb{R}^{d}$, represents a specific temporal path feature that learns the representation of temporal path from subject $s$ to a particular candidate object $o$, where $o\in\mathcal{V}$.
@@
(2) $\textbf{h}^{\prec t+1}_{(s,r)\rightarrow o}=\bigoplus\mathcal{P}_{s\rightarrow o}^{\prec t+1}=\textbf{p}_{1}\oplus\textbf{p}_{2}\oplus\cdot\cdot\cdot\oplus\textbf{p}_{|\mathcal{P}_{s\rightarrow o}^{\prec t+1}|}\Big{|}_{{\rm p}_{k}\in\mathcal{P}_{s\rightarrow o}^{\prec t+1}},$
@@
(3) $\textbf{p}_{k}^{t}=\textbf{w}_{r}(e_{1})\otimes\textbf{w}_{r}(e_{2})\otimes\cdot\cdot\cdot\otimes\textbf{w}_{r}(e_{|{\rm p}_{k}|}),$
@@
Figure 3. An illustrative diagram of the proposed TiPNN model for query-aware temporal path processing.
4.4. Query-aware Temporal Path Processing

### 25 Jan 2024 v3

@@
Abstract
Temporal Knowledge Graph (TKG) is an extension of traditional Knowledge Graph (KG) that incorporates the dimension of time. Reasoning on TKGs is a crucial task that aims to predict future facts based on historical occurrences. The key challenge lies in uncovering structural dependencies within historical subgraphs and temporal patterns. Most existing approaches model TKGs relying on entity modeling, as nodes in the graph play a crucial role in knowledge representation. However, the real-world scenario often involves an extensive number of entities, with new entities emerging over time. This makes it challenging for entity-dependent methods to cope with extensive volumes of entities, and effectively handling newly emerging entities also becomes a significant challenge. Therefore, we propose Temporal Inductive Path Neural Network (TiPNN), which models historical information in an entity-independent perspective. Specifically, TiPNN adopts a unified graph, namely history temporal graph, to comprehensively capture and encapsulate information from history. Subsequently, we utilize the defined query-aware temporal paths on a history temporal graph to model historical path information related to queries for reasoning. Extensive experiments illustrate that the proposed model not only attains significant performance enhancements but also handles inductive settings, while additionally facilitating the provision of reasoning evidence through history temporal graphs.
keywords:
@@ Abstract
[…]
1 Introduction
@@
Knowledge Graphs (KGs) are powerful representations of structured information that capture a vast array of real-world facts. They consist of nodes, which represent entities such as people, places, and events, and edges that define the relationships between these entities. These relationships are typically represented as triples in the form of $(s,r,o)$, comprising a subject, a relation, and an object, such as (Tesla, Founded by, Elon Musk). Temporal Knowledge Graphs (TKGs) is an extension of a traditional KGs that incorporates the dimension of time. In a TKG, these facts are annotated with timestamps or time intervals to indicate when they are or were valid. Different from KGs, each fact in a TKG is represented as a quadruple, adding a timestamp feature, which can be denoted as $(s,r,o,t)$, including a subject, a relation, an object, and a timestamp, such as (Albert Einstein, Born in, Germany, 03/14/1879). The timestamp indicates the specific date or time when this event occurred, providing temporal information to the graph. Such unique representation empowers TKGs to capture the dynamics of multi-relational graphs over time, exhibiting temporal variations. As a result, TKGs have found widespread applications in diverse domains, like social network analysis a272db48509f498facf752cb886353a9 ; doi:10.1177/1525822X16643709 , and event prediction jin2019recurrent ; 10.1145/3443687 ; DBLP:conf/kdd/WangLJLF20 (e.g. disease outbreaks, natural disasters, and political shifts). These applications leverage the temporal information embedded in TKGs to make informed decisions and predictions based on historical knowledge.
@@
Among various tasks, TKG reasoning focuses on inferring missing facts based on known ones and consists of two primary settings: interpolation and extrapolation jin2020Renet . Given the timestamp scope $[0,T]$ in a TKG, where $T$ represents the maximum timestamp, in the interpolation setting, it focuses on the time range from 00 to $T$, i.e., completing missing facts within the range of past timestamps 2020Diachronic . On the other hand, in the extrapolation setting, the reasoning timestamps extend beyond $T$, focusing on predicting the facts in the future li2021temporal ; dong2023adaptive ; 2020Learning . This study particularly emphasizes the extrapolation setting.
@@
Existing literature on TKG reasoning mainly focuses on exploring temporal features and structural information to capture complicated variations and logical connections between entities and relations. For example, RE-NET jin2020Renet and RE-GCN li2021temporal are representative works that leverage RNN-based and GNN-based approaches to learn temporal and structural features of entities in TKGs. xERTE han2020explainable utilizes the entity connection patterns from a constructed inference graph. These methods rely on entity modeling as nodes in the graph play a crucial role in knowledge representation. However, the real-world scenario often involves an extensive number of entities, with new entities emerging over time. This makes it challenging for entity-dependent methods to cope with extensive volumes of entities, and effectively handling newly emerging entities also becomes a significant challenge.
@@
To address this, we previously introduced an innovative entity-independent modeling approach for TKGs with an evolutionary perspective in our previous work, referred to as DaeMon dong2023adaptive . The entity-independent modeling approach brings several benefits: (i) modeling on graph representation is independent of entities, maintaining stable performance even for datasets with numerous entities. Besides, the memory occupation is independent of the number of entity categories. (ii) the entity-independent nature makes it insensitive to entity-specific features. Hence, when new entities emerge in future timestamps, the model can still handle these unseen nodes, showcasing robust generalization and reasoning capabilities. In essence, DaeMon focuses on learning query-aware latent path representations, incorporating a passing strategy between adjacent timestamps to control the path information flow toward later ones without explicitly modeling individual entities. Specifically, DaeMon introduces the concept of temporal paths to represent the logical paths between a query’s subject entity and candidate object entities within the sequence of historical subgraphs. The model employs a path aggregation unit to capture local path information within each subgraph and utilizes path memory units to store aggregated representations of temporal paths along the timeline. Finally, DaeMon adopts a memory passing strategy to facilitate cross-temporal information propagation.
@@
Here, we provide a generalized overview of inference based on temporal paths modeling, as illustrated in Figure 1. For a given inference query, it starts from this query and collects the path between the subject entity and object entity in the historical context. Then, it utilizes a query-aware temporal method to represent the aggregation of temporal paths with temporal information for reasoning. For example, as shown in Figure 1, it iteratively learns the path information between the red triangular and red circular nodes within the historical context. The aggregation of path information corresponding to the query is then utilized to infer the possibility of potential future interactions. Note that, during the learning of path information, the method does not focus on which intermediary nodes (i.e., the blue nodes in Figure 1) are present in the path, but rather employs the relation representations present in the path to express it, showcasing an entity-independent characteristic.
Figure 1: An example of inference based on temporal paths modeling. The red triangular nodes and red circular nodes represent the subject entity and candidate object entity of the query at timestamp $T+1$, respectively.
@@
While DaeMon benefits from this entity-independent modeling approach, it still exhibits certain limitations. First, during the learning of historical information, DaeMon independently performs graph learning on local subgraphs and then connects them through memory passing strategy to obtain the final query-aware path features for future fact reasoning. This process requires graph learning on each subgraph separately, leading to difficulties in modeling complex temporal characteristics and a significant increase in time complexity as the historical length grows. Second, the memory passing strategy is adopted to model temporal features in historical sequences, controlling the information flow from one timestamp to the next to achieve temporal evolution modeling. However, this temporal modeling approach is unidirectional, which not only gives rise to the long-term dependency problem but also neglects relative temporal features between timestamps. Third, DaeMon defines temporal paths as virtually connected paths across the sequence of subgraphs. In practice, it independently learns several edges on each timestamp’s subgraph and then obtains the path feature by fusing the edge information through memory passing strategy. The independent learning approach without real physical paths creates challenges in providing relevant reasoning evidence during the inference stage, which makes it difficult to present interpretable explanations to users.
@@
To overcome the above challenges, here we propose a Temporal Inductive Path Neural Network (TiPNN). Specifically, we introduce the concept of history temporal graphs as a replacement for the original historical subgraph sequences to mitigate the complexity of learning historical information independently on multiple subgraphs. It allows us to model historical information more efficiently by combining it into a single unified graph. Furthermore, we redefine the notion of temporal paths as a logical semantics path between the query entity and candidate entities in history temporal graph and design query-aware temporal path processing to model the path feature on the single unified graph in the entity-independent manner for the final reasoning. It enables us to comprehensively learn both the temporal and structural features present in the history temporal graph. The adoption of the entity-independent approach allows the model to effectively handle the inductive setting on the history temporal graph. By jointly modeling the information within and between different timestamps in a single unified graph, it can capture a more intuitive and holistic view of the temporal reasoning logic, thus facilitating user understanding of the model’s reasoning process and can even provide interpretable explanations for the reasoning results. Our main contributions are as follows:
  • 1.
@@
    We define the concept of temporal paths and propose the Temporal Inductive Path Neural Network (TiPNN) for the task of extrapolated reasoning over TKGs. Our approach models the temporal path features corresponding to reasoning queries by considering temporal information within and between different timestamps, along with the compositional logical rules of historical subgraph sequences.
  • 3.
@@
    Extensive experiments demonstrate that our model outperforms the existing state-of-the-art results. Additionally, we construct datasets and conduct validation experiments for inductive reasoning on TKGs. Furthermore, we present reasoning evidence demonstrations and analytical experiments based on history temporal graph inference.
2 Preliminaries
Table 1: Notations and Descriptions.
@@ 2 Preliminaries
2.1 Background of Temporal Knowledge Graph
@@
2.2 Formulation of Reasoning Task
@@ 3.1 Model Overview
3.2 History Temporal Graph Construction
@@
Figure 2: Example of constructing the history temporal graph $\hat{G}_{t-2:t}$ with $m=3$. For illustrative purposes, we use an undirected graph to demonstrate the construction method and omit the relationship types of edges in the historical subgraph sequence and the constructed history temporal graph. The time labels attached to the edges in the history temporal graph represent the timestamps of the corresponding edges in the historical subgraph sequence.
@@
3.3 Temporal Path Formulation
@@
Definition 1.
Temporal Path is a logical path that aggregates the semantics of all connected paths in history temporal graph between a subject entity and an object entity through arithmetic logical operations.
The aggregated temporal path can be used to simultaneously represent the temporal information and comprehensive path information from the subject entity to the object entity. This enables us to obtain a comprehensive query-aware paired representation for future facts reasoning by aggregating the information from various paths, which is from query subject entity to object candidates in history temporal graph.
Correspondingly, in the previous work dong2023adaptive , DaeMon individually addresses the logical connection patterns within each independent historical subgraph. Additionally, it utilizes a path memory unit to separately model temporal path patterns along the timeline. It’s crucial to note that the temporal paths learned by DaeMon are virtual since they are not acquired on a unified graph, making it impossible for practical visualization on a unified graph. In other words, the temporal path defined in DaeMon represents a latent synthesis of paths across time by the neural network, unlike the explicitly existing path on the history temporal graph in TiPNN.
It is also essential to consider the relation type information from the query while learning paired representations. That is, for a query $(s,r,?,t+1)$, we should also embed the relation type $r$ into the paired representation (i.e., path representation) as well. This ensures that the relation type is incorporated into the learning process, enhancing the overall modeling of the query and enabling more accurate and context-aware inference of missing facts, and the detail will be discussed in Section 3.4. Here we denote the representation of temporal path in history temporal graph $\hat{G}_{\prec t+1}$ as $\textbf{H}^{\prec t+1}_{(s,r)\rightarrow\mathcal{V}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ corresponding to the query $(s,r,?)$ at future timestamp $t+1$, where $|\mathcal{V}|$ is the cardinality of object candidates set and $d$ is the dimension of temporal path embedding. Specifically, $\textbf{H}^{\prec t+1}_{(s,r)\rightarrow\mathcal{V}}$ describes the representations of temporal paths from query subject entity $s$ to all object candidate entities in $\mathcal{V}$, considering the specific query $(s,r,?)$. Each item in $\textbf{H}^{\prec t+1}_{(s,r)\rightarrow\mathcal{V}}$, which is denoted as $\textbf{h}^{\prec t+1}_{(s,r)\rightarrow o}\in\mathbb{R}^{d}$, represents a specific temporal path feature that learns the representation of temporal path from subject $s$ to a particular candidate object $o$, where $o\in\mathcal{V}$.
@@
$\textbf{h}^{\prec t+1}_{(s,r)\rightarrow o}=\bigoplus\mathcal{P}_{s\rightarrow o}^{\prec t+1}=\textbf{p}_{1}\oplus\textbf{p}_{2}\oplus\cdot\cdot\cdot\oplus\textbf{p}_{|\mathcal{P}_{s\rightarrow o}^{\prec t+1}|}\Big{|}_{{\rm p}_{k}\in\mathcal{P}_{s\rightarrow o}^{\prec t+1}},$ (2)
@@
$\textbf{p}_{k}^{t}=\textbf{w}_{r}(e_{1})\otimes\textbf{w}_{r}(e_{2})\otimes\cdot\cdot\cdot\otimes\textbf{w}_{r}(e_{|{\rm p}_{k}|}),$ (3)
@@
3.4 Query-aware Temporal Path Processing